# Adaptive Force-Based Control of Dynamic Legged Locomotion over Uneven Terrain

### v1

@@ IV-A Closed-loop Dynamics
The $L_{1}$ adaptive control is basically designed for trajectory tracking; however, the goal of the balance controller is to compute optimal GRFs. Hence, to integrate the balance controller presented in Sec. II-D into $L_{1}$ adaptive control, we should relate the linear model described in (41) into the closed-loop dynamics.
Let us consider the system state error ($\bm{e}$) according to equation (68) as the state variable. Therefore, the closed-loop error dynamics in state-space form can be represented as follow:
@@
where $\bm{M}$ and $\bm{G}$ are defined in (27). By substituting (82) into the QP problem (II-D), we can obtain the optimal GRFs as the input for the low-level leg controller. The objective of the QP formulation in equation (II-D) is to find a solution that ensures the actual dynamics $\bm{A}\bm{F}$ match the desired dynamics $\bm{b}_{d}$. In general, the QP-based balance controller is capable of achieving the desired control input function outlined in equation (78), thus keeping the error $\bm{e}$ within a certain range. However, if the desired dynamics vector $\bm{b}_{d}$ violates any of the inequality constraints, such as force limits or friction constraints, the controller may yield an optimal solution $\bm{F}^{*}$ that may not completely align with the desired dynamics. With this solution, the optimal dynamic ${\bm{b}_{d}}^{*}$ and $\bm{u}^{*}$ can be written as:
@@
Note that the optimal ground reaction force $\bm{F}^{*}$ serves as the control input for the robot and the variable $\bm{u}^{*}$ acts as an input for the closed-loop dynamic. The closed-loop structure for the robot is depicted in Fig. 3b (the green dashed line).
@@ IV-B Effects of Uncertainty on Dynamic
If we consider uncertainty in the dynamic equation (36) and assume that the matrices $\bm{D}$ and $\bm{H}$ are not accurate, then we need to present the dynamic based on the nominal matrices $\bar{\bm{D}}$, $\bar{\bm{H}}$. The model uncertainty mostly comes from inaccurate values for mass, inertia, and foot position with respect to the center of mass. In addition to that, various terrain (e.g., rough terrain or soft terrain) might have a different impact on the robot, and it is unknown in a practical situation. Therefore, terrain uncertainty should also be considered in the dynamic model. In this section, we solely derive our control equations based on the model uncertainty. In Sec. VI, we will elaborate on how our proposed control system can also consider terrain uncertainty.
There is another parameter involved in the dynamic equation, namely the yaw angle. This angle is obtained through the state estimation, and we assumed that the state estimation has minimal uncertainty. According to the definition of matrices $\bm{D}$ and $\bm{H}$ in (19), the inaccurate value of the dynamic parameter mentioned above reflects on the $\bm{H}$ matrix. Therefore, the dynamic equation in the presence of uncertainty can be represented as:
@@ IV-C Designing Adaptive Controller for Compensating the Uncertainty
By incorporating $L_{1}$ adaptive controller, we want to design a combined controller $\bm{u}=\bm{u}_{1}+\bm{u}_{2}$, where $\bm{u}_{1}$ is the control input to follow the desired trajectory for the nominal model as presented in (78) and $\bm{u}_{2}$ is to compensate the nonlinear model uncertainties $\bm{\theta}$. Therefore, the goal is to adjust the control signal $\bm{u}_{2}$ so that the real model can follow the reference model. For the reference model, we employ a similar linear model described in (41) which, instead of $\bm{M}$, the nominal matrix $\bar{\bm{M}}$ is being used. The diagram of our proposed force-based adaptive control based on a balance controller is presented in Fig. 3b.
@@
Note that the vector of uncertainty $\bm{\theta}$ in equations (88) and (91) are not the same since the state vector of equation (88) is $\bm{X}$ while the state vector of equation (91) is system error $\bm{e}$.
@@
where $\bm{P}\in\mathbb{R}^{12\times 12}$ is a positive definite matrix that is defined according to the stability criteria using the Lyapunov equation. Moreover, the stability proof of the system is provided in the appendix.
@@ V Adaptive Force-based Control using MPC
Model predictive control (MPC) has been widely used across various fields, from finance to robotics. One of MPC’s main advantages is its ability to handle complex systems with multiple inputs and outputs while considering hard control constraints [55]. MPC has also been applied to quadruped robots, providing stable locomotion [1]. Thanks to dynamic prediction in MPC, by using the same control framework, it can achieve different dynamic locomotion gaits. However, MPC’s limitations become evident when dealing with significant uncertainty in the dynamic model. For instance, in the case of a quadruped robot carrying an unknown heavy load, MPC fails to track the desired state trajectory, resulting in unstable behavior and deviation from the desired trajectory, especially with dynamic gaits like bounding. Furthermore, the ability of a robot to traverse soft terrain, where the impact model is unknown, can present a significant challenge. Our proposed approach can tackle this challenge effectively, and we will discuss the details of how it handles the terrain unknown impact model in Sec. VI.
In the previous section Sec. IV, we presented an adaptive force-based control framework based on the balance controller. The balance controller relies on a quadratic program (QP) solver, which is simple to put into practice and well-suited for motions that are slow and safe, like standing and quasi-static walking. Additionally, the balance controller is an instantaneous control technique, meaning it does not predict the robot’s future movement. As a result, the balance controller proves to be ineffective in fast-paced, highly dynamic scenarios. On the other hand, MPC has shown great potential in handling agile motions, even when it comes to underactuated gaits such as bounding.
In this section, we will present a novel control architecture to integrate adaptive control into the MPC framework. By this proposed framework, we can achieve fast and robust locomotion in the presence of uncertainties. This framework can also be extended to accommodate various dynamic gaits, such as trotting and bounding, in legged robots. As we discussed in a previous section, our approach is not restricted to a specific type of adaptive control, but we have chosen to utilize $L_{1}$ adaptive control, which has demonstrated advantages over other adaptive control techniques. The first step in integrating $L_{1}$ adaptive control and MPC is to understand the importance of a reference model and the challenges in synchronizing the real model and reference model. We then present our proposed adaptive MPC, which combines conventional MPC [1] with adaptive control. Finally, we address the challenge of real-time computation while having two MPCs in our control system. We will elaborate on how to adjust the frequency of each control component in an optimized manner to allocate enough computation resources for critical control parts and achieve real-time computation.
@@ V-A Reference Model
Our method aims to design a combined controller based on MPC and $L_{1}$ adaptive control that the real model follows the reference model. In accordance with our previous discussion in Sec. IV-C, the combined controller incorporates a control signal $\bm{u}_{2}$ to account for model uncertainty, as indicated in equation (91). In this section, the auxiliary control signal for this purpose is $\bm{u}_{a}\in\mathbb{R}^{6}$, thus, the uncertain dynamic equation (88) can be rewritten as follow:
@@
We consider a reference model for $L_{1}$ adaptive control that arises from MPC. The MPC method is computationally expensive, but replacing it with other simpler control methods, such as the balance controller while simulating the robot’s performance using dynamic gaits such as bounding is impossible. The reason is that in bounding gait, the robot’s two feet on either the front or rear side touch the ground at each time step, making it challenging to accurately control the height and pitch angle. The MPC approach balances the error in the height and pitch angle and, based on the predicted dynamics of the system in the future, computes the optimal ground reaction forces. As seen in Fig. 4, the center of mass (COM) height oscillates around the desired value. Thus, the underactuated nature of certain gaits like bounding necessitates the use of MPC as the control system for the reference model.
@@
Figure 4: Motion snapshot of the robot with bounding gaits. The quadruped’s center of mass motion (yellow line) cannot be easily predicted with a simple controller. This illustration can represent the importance of using MPC for reference model
@@ V-C Real-time Computation
The main challenge in executing our proposed adaptive MPC framework is ensuring that the computation required is fast enough to be performed in real-time for hardware experiments. If the controller is unable to perform updates at a high frequency, it could result in the robot collapsing during dynamic motion. The control system comprises two MPCs, each with 13 to 19 states predicted over ten horizons. To ensure the robot’s balance and allocate sufficient computation resources to each control component, we have devised a scheme, as depicted in Fig. 3c, to update each control component in an optimized manner.
@@
One of the critical components in our proposed framework is the adaptive MPC, which is responsible for computing the ground reaction force for the robot, as shown in Fig. 3c). Through our experimentation, we have determined that for robust locomotion with dynamic gaits, the optimal update frequency for the adaptive MPC should be 300 Hz. In contrast, the reference MPC, which plays a supporting role in the control system, is less sensitive and runs at a slower rate of 30 Hz. In addition, there is a two-millisecond delay between the running of the adaptive MPC and reference MPC to ensure sufficient computational resources are allocated to each component. This means that the two MPC frameworks do not run simultaneously in our control system.
@@ VI Adaptation to Unknown Impact model
The dynamic formulation presented in Sec. IV and Sec. V considers the presence of model uncertainty in real-world situations. It is assumed that the terrain is hard enough to allow the robot receives the desired force as ground reaction forces on its feet. However, this assumption may not hold if the robot walks on soft or elastic terrain with an unknown impact model, which may not generate the desired force needed for stable locomotion. Some previous studies have included terrain knowledge and contact models in their balancing controllers to address the soft terrain challenge, mainly using a spring-damper model to characterize the soft terrain [58, 59]. Some control frameworks for adapting to soft terrain in real-time have also been developed using iterative learning [60] and whole-body control [61], without prior knowledge about the terrain. This section demonstrates that the proposed method in sections Sec. IV and Sec. V can also handle unknown impact models from terrain, allowing the robot to maintain stability while walking on soft terrains.
@@
which $\bm{F}_{a}$ is the actual ground reaction force exerted on the robot and $\tilde{\bm{F}}_{a}$ is the difference between the desired ground reaction force and actual reaction force. Given that $\tilde{\bm{F}}_{a}$ depends on the tracking error $\bm{e}$ and time, the uncertainty vector arising from the ground reaction force can be incorporated with $\bm{\theta}$. Therefore, we can reformulate equation (131) as follows:
@@ VII Results
In this section, we validate our control approach in simulation and hardware experiments on a Unitree A1 robot. All the hardware experiment’s computation runs on a single PC (Intel i7-6500U, 2.5 GHz, 64-bit). For simulation, the control system is implemented in ROS Noetic with the Gazebo 11 simulator, which provides a high-fidelity simulation of the A1 robot. A video showcasing the results accompanies this paper [^1] https://youtu.be/QmwyysdTk1k.
@@ VII-A1 Walking with significant model uncertainty
The experiment involves the robot walking and rotating in different directions, using both adaptive and non-adaptive controllers while carrying an unknown load. The results of the experiment show that the adaptive controller provides robust locomotion, with excellent tracking error, even when carrying an unknown 5 kg load. On the other hand, the non-adaptive controller results in a considerable error in the COM height and eventually collapses under the weight of just a 3 kg load. The comparative results for the adaptive and non-adaptive controllers are shown in Fig. 6.
@@
Figure 6: Comparing performance of adaptive and non-adaptive controllers. a) Snapshots of the A1 robot with the non-adaptive controller while carrying an unknown 3 kg load and it collapses, b) Snapshots of the A1 robot walking robustly with the adaptive controller while carrying an unknown 5 kg load, c) Comparative plots of the COM height for adaptive and non-adaptive controllers.
@@ VII-A2 Walking on soft terrain
To evaluate the capability of our proposed control method in handling unknown impact models, we conducted an experiment where the robot was made for walking on a double foam, which symbolizes a soft terrain. The performance of both the adaptive and non-adaptive controllers was evaluated and compared. The results are depicted in Fig. 7, which represents the robot’s roll angle. The figure clearly illustrates that the adaptive controller was able to maintain the robot’s balance on the soft terrain, while the non-adaptive controller was unable to do so, leading to the collapse of the robot.
@@ VII-B Running with Multiple Gaits
To demonstrate the superiority of our proposed approach for dynamic gaits, we conducted experiments with the robot running while carrying an unknown load. These experiments were carried out for both the trotting and bounding gaits, with an unknown load of 5 kg and 3 kg, respectively. The results of these experiments are shown in 8. It can be seen from the figure that the tracking of the center of mass height during the bounding gait is more unstable compared to the trotting gait, which is due to the inherent underactuated nature of the bounding gait.

### v2

@@ IV-A Closed-loop Dynamics
The $L_{1}$ adaptive control is designed for trajectory tracking; however, the goal of the balance controller is to compute optimal GRFs. Hence, to integrate the balance controller presented in Sec. II-D into $L_{1}$ adaptive control, we should relate the linear model described in (41) into the closed-loop dynamics.
Let us consider the system state error ($\bm{e}$) according to equation (68) as the state variable. Therefore, the closed-loop error dynamics in state-space form can be represented as follows:
@@
where $\bm{M}$ and $\bm{G}$ are defined in (27). By substituting (82) into the QP problem (II-D), we can obtain the optimal GRFs as the input for the low-level leg controller. The objective of the QP formulation in equation (II-D) is to find a solution that ensures the actual dynamics $\bm{A}\bm{F}$ match the desired dynamics $\bm{b}_{d}$. The QP-based balance controller can generally achieve the desired control input function outlined in equation (78), thus keeping the error $\bm{e}$ within a certain range. However, if the desired dynamics vector $\bm{b}_{d}$ violates any of the inequality constraints, such as force limits or friction constraints, the controller may yield an optimal solution $\bm{F}^{*}$ that may not completely align with the desired dynamics. With this solution, the optimal dynamic ${\bm{b}_{d}}^{*}$ and $\bm{u}^{*}$ can be written as:
@@
Note that the optimal ground reaction force $\bm{F}^{*}$ serves as the control input for the robot, and the variable $\bm{u}^{*}$ acts as an input for the closed-loop dynamic. The closed-loop structure for the robot is depicted in Fig. 3b (the green dashed line).
@@ IV-B Effects of Uncertainty on Dynamic
If we consider uncertainty in the dynamic equation (36) and assume that the matrices $\bm{D}$ and $\bm{H}$ are not accurate, then we need to present the dynamic based on the nominal matrices $\bar{\bm{D}}$, $\bar{\bm{H}}$. The model uncertainty mostly comes from inaccurate values for mass, inertia, and foot position with respect to the center of mass. In addition, various terrains (e.g., rough terrain or soft terrain) might have different impacts on the robot, which is unknown in a practical situation. Therefore, terrain uncertainty should also be considered in the dynamic model. In this section, we solely derive our control equations based on the model uncertainty. In Sec. VI, we will elaborate on how our proposed control system can also consider terrain uncertainty.
Another parameter is involved in the dynamic equation, namely the yaw angle. This angle is obtained through the state estimation, and we assumed that the state estimation has minimal uncertainty. According to the definition of matrices $\bm{D}$ and $\bm{H}$ in (19), the inaccurate value of the dynamic parameter mentioned above reflects on the $\bm{H}$ matrix. Therefore, the dynamic equation in the presence of uncertainty can be represented as:
@@ IV-C Designing Adaptive Controller for Compensating the Uncertainty
By incorporating $L_{1}$ adaptive controller, we want to design a combined controller $\bm{u}=\bm{u}_{1}+\bm{u}_{2}$, where $\bm{u}_{1}$ is the control input to follow the desired trajectory for the nominal model as presented in (78) and $\bm{u}_{2}$ is to compensate the nonlinear model uncertainties $\bm{\theta}$. Therefore, the goal is to adjust the control signal $\bm{u}_{2}$ so that the real model can follow the reference model. For the reference model, we employ a similar linear model described in (41) which, instead of $\bm{M}$, the nominal matrix $\bar{\bm{M}}$ is being used. The proposed force-based adaptive control diagram based on a balance controller is presented in Fig. 3b.
@@
Note that the vector of uncertainty $\bm{\theta}$ in equations (88) and (91) are not the same since the state vector of equation (88) is $\bm{X}$. In contrast, the state vector of equation (91) is system error $\bm{e}$.
@@
where $\bm{P}\in\mathbb{R}^{12\times 12}$ is a positive definite matrix that is defined according to the stability criteria using the Lyapunov equation. Moreover, the system’s stability proof is provided in the appendix.
@@ V Adaptive Force-based Control using MPC
Model predictive control (MPC) has been widely used across various fields, from finance to robotics. One of MPC’s main advantages is its ability to handle complex systems with multiple inputs and outputs while considering hard control constraints [64]. MPC has also been applied to quadruped robots, providing stable locomotion [1]. Thanks to dynamic prediction in MPC, using the same control framework can achieve different dynamic locomotion gaits. However, MPC’s limitations become evident when dealing with significant uncertainty in the dynamic model. For instance, in the case of a quadruped robot carrying an unknown heavy load, MPC fails to track the desired state trajectory, resulting in unstable behavior and deviation from the desired trajectory, especially with dynamic gaits like bounding. Furthermore, the ability of a robot to traverse soft terrain where the impact model is unknown can present a significant challenge. Our proposed approach can tackle this challenge effectively, and we will discuss how it handles the terrain unknown impact model in Sec. VI.
In the previous section Sec. IV, we presented an adaptive force-based control framework based on the balance controller. The balance controller relies on a quadratic program (QP) solver, which is simple to put into practice and well-suited for slow and safe motions like standing and quasi-static walking. Additionally, the balance controller is an instantaneous control technique, meaning it does not predict the robot’s future movement. As a result, the balance controller proves to be ineffective in fast-paced, highly dynamic scenarios. On the other hand, MPC has shown great potential in handling agile motions, even when it comes to underactuated gaits such as bounding.
In this section, we will present a novel control architecture to integrate adaptive control into the MPC framework. By this proposed framework, we can achieve fast and robust locomotion in the presence of uncertainties. This framework can also be extended to accommodate various dynamic gaits in legged robots, such as trotting and bounding. As discussed in a previous section, our approach is not restricted to a specific type of adaptive control. Still, we have chosen to utilize $L_{1}$ adaptive control, which has demonstrated advantages over other adaptive control techniques. The first step in integrating $L_{1}$ adaptive control and MPC is understanding the importance of a reference model and the challenges in synchronizing the real and reference models. We then present our proposed adaptive MPC, which combines conventional MPC [1] with adaptive control. Finally, we address the challenge of real-time computation while having two MPCs in our control system. We will elaborate on how to adjust the frequency of each control component in an optimized manner to allocate enough computation resources for critical control parts and achieve real-time computation.
@@ V-A Reference Model
Our method aims to design a combined controller based on MPC and $L_{1}$ adaptive control that the real model follows the reference model. In accordance with our previous discussion in Sec. IV-C, the combined controller incorporates a control signal $\bm{u}_{2}$ to account for model uncertainty, as indicated in equation (91). In this section, the auxiliary control signal for this purpose is $\bm{u}_{a}\in\mathbb{R}^{6}$; thus, the uncertain dynamic equation (88) can be rewritten as follow:
@@
We consider a reference model for $L_{1}$ adaptive control that arises from MPC. The MPC method is computationally expensive, but replacing it with other simpler control methods, such as the balance controller, while simulating the robot’s performance using dynamic gaits such as bounding is impossible. The reason is that in bounding gait, the robot’s two feet on either the front or rear side touch the ground at each time step, making it challenging to accurately control the height and pitch angle. The MPC approach balances the error in the height and pitch angle and, based on the predicted dynamics of the system in the future, computes the optimal ground reaction forces. As seen in Fig. 4, the center of mass (COM) height oscillates around the desired value. Thus, the underactuated nature of certain gaits, such as bounding, necessitates the use of MPC as the control system for the reference model.
@@
Figure 4: Motion snapshot of the robot with bounding gaits. A simple controller cannot easily predict The quadruped’s center of mass motion (yellow line). This illustration can represent the importance of using MPC for reference model
@@ V-C Real-time Computation
The main challenge in executing our proposed adaptive MPC framework is ensuring that the computation required is fast enough for hardware experiments in real time. If the controller is unable to perform updates at a high frequency, it could result in the robot collapsing during dynamic motion. The control system comprises two MPCs, each with 13 to 19 states predicted over ten horizons. To ensure the robot’s balance and allocate sufficient computation resources to each control component, we have devised a scheme, as depicted in Fig. 3c, to update each control component in an optimized manner.
@@
One of the critical components in our proposed framework is the adaptive MPC, which is responsible for computing the ground reaction force for the robot, as shown in Fig. 3c. Through our experimentation, we have determined that for robust locomotion with dynamic gaits, the optimal update frequency for the adaptive MPC should be 300 Hz. In contrast, the reference MPC, which plays a supporting role in the control system, is less sensitive and runs at a slower rate of 30 Hz. In addition, there is a two-millisecond delay between the running of the adaptive MPC and reference MPC to ensure sufficient computational resources are allocated to each component. This means the two MPC frameworks do not run simultaneously in our control system.
@@ VI Adaptation to Unknown Impact model
The dynamic formulation presented in Sec. IV and Sec. V considers the presence of model uncertainty in real-world situations. It is assumed that the terrain is hard enough to allow the robot to receive the desired force as ground reaction forces on its feet. However, this assumption may not hold if the robot walks on soft or elastic terrain with an unknown impact model, which may not generate the desired force needed for stable locomotion. Some previous studies have included terrain knowledge and contact models in their balancing controllers to address the soft terrain challenge, mainly using a spring-damper model to characterize the soft terrain [67, 68]. Some control frameworks for adapting to soft terrain in real-time have also been developed using iterative learning [69] and whole-body control [70], without prior knowledge about the terrain. This section demonstrates that the proposed method in sections Sec. IV and Sec. V can also handle unknown impact models from terrain, allowing the robot to maintain stability while walking on soft terrains.
Equation (70), representing the force-to-torque mapping, holds under the condition that the movement of each leg can be considered negligible. This assumption is reasonable for the stance leg on solid ground. However, when dealing with soft terrain, this mapping is not accurate. The dynamic equation for each leg is expressed as follows:
$\bm{\tau}_{stance,i}=\bm{M}_{i}(\bm{q}_{i})\ddot{\bm{q}_{i}}+\bm{n}_{i}(\bm{q}_{i},\dot{\bm{q}_{i}})-{\bm{J}(\bm{q}_{i})}^{T}\bm{R}^{T}\bm{F}_{i}$ (131)
where $\bm{M}_{i}(\bm{q}_{i})$ is the inertia matrix, $\bm{n}_{i}(\bm{q}_{i},\dot{\bm{q}_{i}})$ is the nonlinear term, and $\ddot{\bm{q}_{i}}$ is joints acceleration for the $i$-th leg. Remember that assumption 1 in Sec. II-C was considered for the rigid body dynamic equation of the robot. However, when addressing the dynamic equation of each leg, it is not valid to neglect the inertia associated with each leg. Furthermore, calculating the joint acceleration poses challenges due to considerable noise, making the complete computation of (131) difficult. This is where adaptive controllers prove to be beneficial.
@@
where $\bm{F}_{a}$ is the actual ground reaction force applied to the robot and $\tilde{\bm{F}}_{a}$ is the difference between the desired ground reaction force and actual reaction force. Therefore, we can reformulate equation (132) as follows:
@@ VII Results
In this section, we validate our control approach in simulation and hardware experiments on a Unitree A1 robot. All the hardware experiments’ computations run on a single PC (Intel i7-6500U, 2.5 GHz, 64-bit). For simulation, the control system is implemented using ROS Noetic in the Gazebo 11 simulator, which provides a high-fidelity simulation of the A1 robot. A video showcasing the results accompanies this paper [^1] https://youtu.be/5t1mSh0q3lk.
@@ VII-A1 Walking with significant model uncertainty
The experiment involves the robot walking and rotating in different directions, using both adaptive and non-adaptive controllers while carrying an unknown load. To ensure a precise comparison, we establish a unit test comprising velocity and rotation commands. This test is then executed for both adaptive and non-adaptive controllers. The experiment results show that the adaptive controller provides robust locomotion, with excellent tracking error, even when carrying an unknown 5 kg load. On the other hand, the non-adaptive controller results in a considerable error in the COM height and eventually collapses under the weight of just a 3 kg load. The comparative results for the adaptive and non-adaptive controllers are shown in Fig. 6.
@@
Figure 6: Comparing performance of adaptive and non-adaptive controllers. a) Snapshots of the A1 robot with the non-adaptive controller while carrying an unknown 3 kg load and collapses, b) Snapshots of the A1 robot walking robustly with the adaptive controller while carrying an unknown 5 kg load. c) Comparative plots of the COM height, d) velocity command tracking in the x-direction, and e) y-direction for adaptive and non-adaptive controllers. The plots for the non-adaptive controller do not persist until the completion of the unit test, as the robot experiences a collapse during testing.
@@ VII-A2 Walking on soft terrain
To evaluate the capability of our proposed control method in handling unknown impact models, we conducted an experiment where the robot was made for walking on a double foam, which symbolizes a soft terrain. The performance of both the adaptive and non-adaptive controllers was evaluated and compared. The results are depicted in Fig. 7, representing the robot’s roll angle. The figure clearly illustrates that the adaptive controller could maintain the robot’s balance on the soft terrain. In contrast, the non-adaptive controller could not do so, leading to the collapse of the robot.
@@ VII-B Running with Multiple Gaits
To demonstrate the superiority of our proposed approach for dynamic gaits, we conducted experiments with the robot running while carrying an unknown load. These experiments were carried out for both the trotting and bounding gaits, with an unknown load of 5 kg and 3 kg, respectively. The results of these experiments are shown in Fig. 8. It can be seen from the figure that the tracking of the center of mass height during the bounding gait is more unstable compared to the trotting gait, which is due to the inherent underactuated nature of the bounding gait.